OpenStudy (anonymous):

if the walls in a room are 955 square feet in area, and a gallon of paint covers 15 square yards, how many gallons of paint are needed for the room? (3 ft=1 yd)

OpenStudy (stormfire1):

1 square yard = 9 square feet so 15 * 9 = 135 square feet which can be painted with one gallon of paint. Divide 955 by 135 to get ~7.074 gallons.

OpenStudy (anonymous):

you figure out how to do these by getting the units to cancel:\[955 ft^2 \times \frac{1 yd}{3 ft} \times \frac{1 yd}{3 ft} \times \frac{1 gal}{15 yd^2} = 7.074 gal\](feet and yards cancel out)